Grade 12th pass10 grade maths

4. The wheels of a car are of diameter 80 cm each. How many complete revolutions does each wheel make in 10 minutes when the car is travelling at a speed of 66 km per hour?

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
The radius of car’s wheel = 80/2 = 40 cm (as D = 80 cm) So, the circumference of wheels = 2πr = 80 π cm Now, in one revolution, the distance covered = circumference of the wheel = 80 π cm It is given that the distance covered by the car in 1 hr = 66km Converting km into cm we get, Distance covered by the car in 1hr = (66×105) cm In 10 minutes, the distance covered will be = (66×105×10)/60 = 1100000 cm/s ∴ Distance covered by car = 11×105 cm Now, the no. of revolutions of the wheels = (Distance covered by the car/Circumference of the wheels) =( 11×105)/80 π = 4375.